# Graph Signal Processing for Global Stock Market Volatility Forecasting

### v1

@@ 4.5 Dynamic Modeling
As mentioned in the Literature Review section, the pattern of interconnectedness among stock markets can change over time. However, there is a paucity of work on incorporating the dynamic volatility spillover effect into RV forecasting models. The proposed GSPHAR model offers potential solutions to handle the changing correlations in the global stock market. The GSPHAR model merges the volatility spillover effect into the RV data through GFT, in which the volatility network is precomputed and independent of the model training process. As long as the volatility network varies across different inputs, the model can become sensitive to the changes in volatility dynamics and adapt to the latest volatility patterns. However, how to change the volatility network with respect to the input RV data is an open question.
@@
where $\rho$ is a non-negative hyperparameter to balance the influence of mid-term historical RVs and long-term historical RVs when calculating $\mathbf{A}$. The outcome matrix $\mathbf{A}$ is leveraged to calculate the magnet Laplacian matrix, which is used in the GSPHAR model. The correlation coefficient matrices $|\mathbf{A}^{\text{P}}_{t-5:t-1}|$ and $|\mathbf{A}^{\text{P}}_{t-22:t-1}|$ can be regarded as a modification or adjustment applied to the magnitude of the overall volatility interconnection pattern $\mathbf{A}^{\text{DY}}$ such that the overall volatility network is adjusted towards the more recent RV data. A GSPHAR model that incorporates such dynamic designs is referred to as the dynamic-GSPHAR (d-GSPHAR) model.
@@ 5 Experiments
This section empirically evaluates the performance of the proposed v-GSPHAR, GSPHAR and d-GSPHAR models, using multivariate square root RV time series from $24$ stock market indices. The experiments are designed to compare these proposed models against baseline models to evaluate the validity and effectiveness of the GSP-based RV forecasting methods and to uncover new empirical insights into RV dynamics. The usefulness of the convolution filters in the GSPHAR model is also investigated. Furthermore, the d-GSPHAR model is compared with the GSPHAR model to determine the effect of incorporating volatility dynamics into the proposed framework.
@@ 5.1 Model Evaluation Tools
Various evaluation criteria are applied to evaluate the effectiveness of the proposed models. The evaluation process is conducted on an index-wise basis, as different markets operate under different economic conditions and investor behaviors. Especially, the accuracy of out-of-sample RV forecasts is measured through the commonly used Mean Absolute Error (MAE, calculation detail to be shown in the following Section 5.3). In addition, the Model Confidence Set (MCS) test (Hansen et al., 2011) and the Diebold-Mariano (DM) test (Diebold and Mariano, 2002) are applied to compare the accuracy of out-of-sample forecasts between the proposed models and baseline models to assess whether the improvement is significant or not. Similar to the input settings of the HAR models, the out-of-sample forecasting windows are set as $H=1,5,22$, corresponding to the short-term (daily), mid-term (weekly) and long-term (monthly) forecasting.
@@
  • Software:
    • Python 3.10.14,
    • NumPy 1.26.4,
    • PyTorch 1.13.1+cu116.
  • Hardware:
    • Operating system: Windows 11,
    • CPU: 13th Gen Intel(R) Core(TM) i9-13900HX,
    • GPU: NVIDIA GeForce RTX 4080 Laptop GPU.
@@ 5.2 Benchmark Datasets and Baseline models
The benchmark dataset consists of the square root RV data of $24$ stock market indices. Among them, some indices are commonly used indices in previous RV forecast research (e.g., Liang et al. (2020)). The daily RV data is produced by Equation (1) based on the $5$-minute high-frequency returns. Given that multivariate RV forecast models are trained on RV observations on common trading days of the included stock market indices, the selection of these $24$ indices is intentionally focused on markets with more overlapping trading schedules. This approach ensures that the RV time series can be more continuous with more synchronized trading days across different markets, thus enhancing the reliability and comparability of the multivariate RV forecasts produced by different models. Besides, the whole dataset is scaled up by multiplying $100$.
A descriptive analysis of the scaled square-rooted data for each stock market index is reported in Table 11 in Appendix A. The Augmented Dickey-Fuller (ADF) test (MacKinnon, 1994; Cheung and Lai, 1995) is conducted and results reject the existence of the unit root in each time series at the $5\%$ significance level. This indicates that the time series data are stationary. As illustrated in the Volatility Spillover Graph Formulation section, the DY framework is applied to measure the net pairwise volatility spillover effect between different stock market indices. The net volatility spillover graph of the $24$ stock market indices calculated following Equation (26) is presented below. The nodes represent different indices and the numbers on the edges measure the net pairwise volatility spillover effect. Larger values indicate stronger volatility spillover effects while smaller values reflect weaker volatility transmission.
Figure 3: Net volatility spillover graph
The baseline models include the HAR, VHAR, HAR-KS and GNNHAR presented in the Literature Review section. Although the exact construction of some baseline models is not available, they are formulated according to the corresponding research papers. The relevant dataset and the code of the proposed models can be found at https://github.com/MikeZChi/GSPHAR.git.
@@ 5.3 Experiment Results
where $\text{MAE}_{i}$ is the MAE of the $i_{\text{th}}$ stock market index and $T_{\text{out}}$ denotes the duration of the out-of-sample dataset. $\widehat{\mathbf{v}}_{i,t}$ refers to the square root RV forecasting value of the $i_{\text{th}}$ index at time $t$ and $\mathbf{v}_{i,t}$ is the true square root RV value of the $i_{\text{th}}$ index at time $t$. The index-wise MAE results produced by each model with different forecast window $H$ are presented in Tables 1, 2 and 3.
@@
$H=1$ HAR VHAR HAR-KS GNNHAR v-GSPHAR GSPHAR
@@
$H=5$ HAR VHAR HAR-KS GNNHAR v-GSPHAR GSPHAR
@@
$H=22$ HAR VHAR HAR-KS GNNHAR v-GSPHAR GSPHAR
@@
According to the MAE results in Tables 1, 2 and 3, it is clear that the GSPHAR model has improved RV forecasting results over all three forecast horizons. The GSPHAR model achieves the minimum MAE scores for $23$, $22$ and $21$ stock market indices in the short-term, mid-term and long-term RV forecasting tasks, respectively. Even when it does not achieve the best performance, the GSPHAR model still produces the second smallest MAE scores. These imply the consistent superiority of the GSPHAR model in different forecasting horizons. Besides, compared to the linear HAR family of models, including the HAR model, the VHAR and the HAR-KS model, the v-GSPHAR model performs better (achieves a smaller MAE score more often) in all forecasting tasks. This indicates that even the basic GSP techniques are more effective than the residual learning method of the baseline HAR models in integrating the volatility spillover effect. However, the v-GSPHAR model is inferior to the GSPHAR model and the GNNHAR model, especially in mid-term and long-term RV forecasting tasks. This implies that the linear expressiveness of the v-GSPHAR model could be insufficient in capturing the potential nonlinear volatility patterns, especially for the mid-term and long-term forecasts.
@@
Table 4: The MCS test results table of the short-term forecasting task with bold numbers representing that the corresponding models are included in the MCS at $5\%$ level.
$H=1$ HAR VHAR HAR-KS GNNHAR v-GSPHAR GSPHAR
FCHI $\mathbf{0.066}$ 00 $120.0$ $\mathbf{1}$ $\mathbf{0.957}$
AEX 00 00 00 00 $\mathbf{0.522}$ $\mathbf{1}$
BFX $040.0$ $\mathbf{0.526}$ 00 $\mathbf{0.615}$ $\mathbf{1}$
STOXX50E $150.0$ $\mathbf{0.175}$ 00 $\mathbf{0.195}$ $\mathbf{1}$
IBEX $030.0$ 00 00 $\mathbf{0.263}$ $\mathbf{1}$
GDAXI $010.0$ 00 00 $0.001$ $\mathbf{1}$
AORD $0.019$ $\mathbf{0.108}$ $\mathbf{0.781}$ $0.004$ $\mathbf{0.561}$ $\mathbf{1}$
FTSE $0.002$ $\mathbf{0.144}$ $\mathbf{0.262}$ $0.001$ $\mathbf{0.837}$ $\mathbf{1}$
MXX $\mathbf{0.559}$ $0.015$ $\mathbf{0.155}$ $\mathbf{0.113}$ $\mathbf{0.559}$ $\mathbf{1}$
IXIC $\mathbf{0.056}$ 00 $0.009$ $0.007$ $0.015$ $\mathbf{1}$
SSMI $150.0$ 00 00 $\mathbf{0.111}$ $\mathbf{1}$
SPX $040.0$ 00 $0.001$ $0.004$ $\mathbf{1}$
RUT $0.040$ $\mathbf{0.761}$ $\mathbf{0.948}$ $0.012$ $\mathbf{0.827}$ $\mathbf{1}$
DJI $010.0$ $0.001$ $0.001$ $0.020$ $\mathbf{1}$
BSESN 00 00 00 00 $0.002$ $\mathbf{1}$
NSEI 00 00 00 00 $0.005$ $\mathbf{1}$
KS11 00 00 00 00 00 $\mathbf{1}$
BVSP $010.0$ 00 00 $\mathbf{0.096}$ $\mathbf{1}$
HSI $0.010$ $0.001$ $0.010$ $0.010$ $\mathbf{0.077}$ $\mathbf{1}$
KSE 00 00 00 00 00 $\mathbf{1}$
N225 00 $010.0$ 00 $0.001$ $\mathbf{1}$
SSEC 00 00 00 00 00 $\mathbf{1}$
OSEAX 00 00 $\mathbf{0.380}$ 00 00 $\mathbf{1}$
GSPTSE $160.0$ 00 $0.016$ $0.035$ $\mathbf{1}$
@@
Table 5: The MCS test results table of the mid-term forecasting task with bold numbers representing that the corresponding models are included in the MCS at $5\%$ level.
$H=5$ HAR VHAR HAR-KS GNNHAR v-GSPHAR GSPHAR
FCHI 00 00 00 00 00 $\mathbf{1}$
AEX $040.0$ $010.0$ $0.004$ $\mathbf{1}$
BFX $120.0$ $0.012$ $0.001$ $0.012$ $\mathbf{1}$
STOXX50E 00 00 00 00 00 $\mathbf{1}$
IBEX 00 00 00 00 00 $\mathbf{1}$
GDAXI 00 00 00 00 00 $\mathbf{1}$
AORD $\mathbf{0.173}$ $0.009$ $\mathbf{1}$ $\mathbf{0.382}$ $\mathbf{0.284}$ $\mathbf{0.713}$
FTSE $280.0$ $\mathbf{0.097}$ $\mathbf{0.329}$ $\mathbf{0.05}$ $\mathbf{1}$
MXX $\mathbf{0.484}$ 00 $\mathbf{0.484}$ 00 $\mathbf{0.484}$ $\mathbf{1}$
IXIC $0.044$ $0.001$ $\mathbf{0.094}$ $\mathbf{0.277}$ $\mathbf{0.094}$ $\mathbf{1}$
SSMI $050.0$ $0.001$ $0.002$ $0.005$ $\mathbf{1}$
SPX $250.0$ $0.014$ $\mathbf{0.795}$ $0.008$ $\mathbf{1}$
RUT $\mathbf{0.187}$ $0.011$ $\mathbf{0.703}$ $\mathbf{0.566}$ $\mathbf{0.765}$ $\mathbf{1}$
DJI $250.0$ $0.046$ $\mathbf{0.587}$ $0.046$ $\mathbf{1}$
BSESN 00 00 00 00 00 $\mathbf{1}$
NSEI 00 00 00 00 00 $\mathbf{1}$
KS11 $0.004$ $0.005$ $040.0$ $0.005$ $\mathbf{1}$
BVSP $400.0$ $\mathbf{0.059}$ $\mathbf{1}$ $\mathbf{0.105}$ $\mathbf{0.767}$
HSI $190.0$ $190.0$ $\mathbf{0.056}$ $\mathbf{1}$
KSE $0.001$ $0.006$ $090.0$ $0.002$ $\mathbf{1}$
N225 $010.0$ $0.001$ $\mathbf{0.555}$ $0.001$ $\mathbf{1}$
SSEC 00 00 00 00 00 $\mathbf{1}$
OSEAX 00 00 $0.001$ $\mathbf{0.606}$ $0.004$ $\mathbf{1}$
GSPTSE $080.0$ $0.001$ $0.008$ $0.008$ $\mathbf{1}$
@@
Table 6: The MCS test results table of the long-term forecasting task with bold numbers representing that the corresponding models are included in the MCS at $5\%$ level.
$H=22$ HAR VHAR HAR-KS GNNHAR v-GSPHAR GSPHAR
FCHI $090.0$ 00 $\mathbf{0.051}$ 00 $\mathbf{1}$
AEX $040.0$ 00 $0.021$ $0.001$ $\mathbf{1}$
BFX $050.0$ 00 $0.005$ $\mathbf{0.473}$ $\mathbf{1}$
STOXX50E $010.0$ 00 $0.007$ $0.001$ $\mathbf{1}$
IBEX $070.0$ 00 $070.0$ $\mathbf{1}$
GDAXI 00 00 00 $080.0$ $\mathbf{1}$
AORD $0.044$ $0.044$ $0.044$ $0.044$ $0.044$ $\mathbf{1}$
FTSE $070.0$ $0.001$ $0.013$ $0.013$ $\mathbf{1}$
MXX $\mathbf{0.594}$ $0.001$ $\mathbf{0.509}$ 00 $\mathbf{0.130}$ $\mathbf{1}$
IXIC $0.050$ $0.020$ $0.050$ $0.050$ $0.050$ $\mathbf{1}$
SSMI $030.0$ 00 $\mathbf{1}$ $0.003$ $\mathbf{0.887}$
SPX $100.0$ $0.002$ $\mathbf{0.200}$ $0.015$ $\mathbf{1}$
RUT $\mathbf{0.089}$ $0.005$ $0.036$ $0.013$ $\mathbf{0.167}$ $\mathbf{1}$
DJI $040.0$ 00 $\mathbf{0.128}$ $0.015$ $\mathbf{1}$
BSESN 00 00 00 $\mathbf{0.389}$ 00 $\mathbf{1}$
NSEI 00 00 00 $\mathbf{0.39}$ 00 $\mathbf{1}$
KS11 $040.0$ $0.008$ $0.021$ $0.004$ $\mathbf{1}$
BVSP $\mathbf{0.111}$ 00 $\mathbf{0.055}$ $\mathbf{0.111}$ $\mathbf{0.111}$ $\mathbf{1}$
HSI $\mathbf{0.271}$ 00 $0.002$ $\mathbf{0.106}$ $\mathbf{0.271}$ $\mathbf{1}$
KSE $190.0$ $\mathbf{0.201}$ $\mathbf{0.218}$ $0.035$ $\mathbf{1}$
N225 00 $030.0$ $\mathbf{0.175}$ $0.024$ $\mathbf{1}$
SSEC 00 00 00 $\mathbf{1}$ 00 $\mathbf{0.570}$
OSEAX $040.0$ 00 $0.008$ $0.008$ $\mathbf{1}$
GSPTSE $030.0$ 00 $\mathbf{1}$ 00 $\mathbf{0.920}$
@@
The MAE results and the MCS test demonstrate that the GSPHAR model improved the RV forecasting performance compared to the other tested models. To further evaluate the RV forecasting results of the GSPHAR model across different forecast horizons, the one-sided DM test is applied to perform pairwise comparisons against the competing models.
Suppose the two candidate models are model 00 and model $1$. Their absolute forecast error series for the $i_{\text{th}}$ stock market index can be represented as $\{e^{(0)}_{i,t}\}_{t=1}^{T}$ and $\{e^{(1)}_{i,t}\}_{t=1}^{T}$. The distance between these two series of absolute forecast errors is measured as $d_{t}=e_{0,t}-e_{1,t}$. The test statistic $D$ of the DM test is calculated as follows:
$D=V_{d}^{-\frac{1}{2}}\overline{d},$ (38)
where $V_{d}$ is the autocorrelation-adjusted variance estimate of $\{d_{t}\}_{t=1}^{T}$ and $\overline{d}$ is the mean distance between the two absolute forecast error series: $\overline{d}=\frac{1}{T}\sum_{t=1}^{T}d_{t}$. According to Equation (38), given that $V_{d}$ is non-negative, $D$ and $\overline{d}$ are both positive, negative or equal to 00. A positive $D$ indicates that, on average, $\{e_{1,t}\}_{t=1}^{T}$ is less than $\{e^{(0)}_{i,t}\}_{t=1}^{T}$ and model $1$ has higher forecast accuracy than model 00. The null hypothesis of the one-sided DM test is that model 00 and model $1$ are indifferent or model $1$ is inferior to model 00 in terms of forecast accuracy. When $D$ is sufficiently large, the DM test can reject the null hypothesis and conclude that model $1$ provides a significantly more accurate forecast than model 00. The results of the DM test are presented in Tables 7, 8 and 9. Here, for each DM test model $1$ represents the GSPHAR model, and model 00 represents other models.
Table 7: The DM test results table of the short-term forecasting task with bold numbers represents that the GSPHAR is significantly better than the competing model at $5\%$ level.

### v2

@@ 3.2 Dynamic Modeling
The pattern of the volatility spillover effect can vary across time because market conditions, economic factors and investor behavior continuously change, which reflects different periods of strengthened or weakened correlations (Diebold and Yilmaz, 2012; Antonakakis et al., 2018; Gong et al., 2021). However, there is a paucity of work on incorporating the dynamic volatility spillover effect into RV forecasting models. The proposed GSPHAR model offers potential solutions to handle the changing correlations in the global stock market. It merges the volatility spillover effect into the RV data through GFT, in which the volatility network is precomputed and independent of the model training process. As long as the volatility network varies across different inputs, the model can become sensitive to the changes in volatility dynamics and adapt to the latest volatility patterns. However, how to change the volatility network with respect to the input RV data is an open question.
@@
where $\rho$ is a non-negative hyperparameter to balance the influence of mid-term historical RVs and long-term historical RVs when calculating $\mathbf{A}$. The outcome matrix $\mathbf{A}$ is leveraged to calculate the magnetic Laplacian matrix, which is used in the GSPHAR model. The correlation coefficient matrices $|\mathbf{A}^{\text{P}}_{t-5:t-1}|$ and $|\mathbf{A}^{\text{P}}_{t-22:t-1}|$ can be regarded as a modification or adjustment applied to the magnitude of the overall volatility interconnection pattern $\mathbf{A}^{\text{DY}}$ such that the overall volatility network is adjusted towards more recent RV data. A GSPHAR model that incorporates such dynamic designs is referred to as the dynamic-GSPHAR (d-GSPHAR) model.
@@ 4 Experiments
This section empirically evaluates the performance of the proposed GSPHAR and d-GSPHAR models, using square root RV time series from $24$ stock market indices. The experiments are designed to compare these proposed models against baseline models to evaluate the validity and effectiveness of the GSPHAR model and to uncover new empirical insights into RV dynamics. The usefulness of the convolution filters in the GSPHAR model is also investigated. Furthermore, the d-GSPHAR model is compared with the GSPHAR model to determine the effect of incorporating volatility dynamics into the proposed framework.
@@ 4.1 Model Evaluation Tools
Various evaluation criteria are applied to evaluate the effectiveness of the proposed models. The evaluation process is conducted on an index-wise basis, as different markets operate under different economic conditions and investor behaviors. Especially, the accuracy of out-of-sample RV forecasts is measured through the commonly used Mean Absolute Error (MAE, calculation detail to be shown in the following Section 4.3). In addition, the Model Confidence Set (MCS) test (Hansen et al., 2011) and the Diebold-Mariano (DM) test (Diebold and Mariano, 1995) are applied to compare the accuracy of out-of-sample forecasts between the proposed models and baseline models to assess whether the improvement is significant or not. Similar to the input settings of the HAR models, the out-of-sample forecasting windows are set as $H=1,5,22$, corresponding to the short-term (daily), mid-term (weekly) and long-term (monthly) forecasting.
@@
  • Software: Python 3.10.14; NumPy 1.26.4; Statsmodels 0.14.2; PyTorch 1.13.1+cu116.
  • Hardware: Operating system: Windows 11; CPU: 13th Gen Intel(R) Core(TM) i9-13900HX; GPU: NVIDIA GeForce RTX 4080 Laptop GPU.
@@ 4.2 Benchmark Datasets and Baseline models
The benchmark dataset consists of the square root RV data of $24$ stock market indices with around $3500$ common trading days from May 2002 to June 2022. Among them, some indices are commonly used indices in previous RV forecast research e.g., Liang et al. (2020). The daily RV data is produced based on the $5$-minute intraday high frequency returns. Given that RV forecast models are trained on RV observations on common trading days of the included stock market indices, the selection of these $24$ indices is intentionally focused on markets with more overlapping trading schedules. This approach ensures that the RV time series can be more continuous with more synchronized trading days across different markets, thus improving the data usage efficiency and enhancing the reliability and comparability of the RV forecasts produced by different models. The first $70\%$ of data (from May 2002 to Feb 2016) are the in-sample data, whereas the last $30\%$ of data (from March 2016 to June 2022) are the out-sample data. Besides, the square root RV is scaled up by multiplying $100$.
A descriptive analysis of the scaled square-rooted data for each stock market index is reported in Table 11 in Appendix A. The Augmented Dickey-Fuller (ADF) test (MacKinnon, 1994; Cheung and Lai, 1995) is conducted and results reject the existence of the unit root in each time series at the $5\%$ significance level. This indicates that the time series data are stationary. As illustrated in the Volatility Spillover Graph Formulation section, the DY framework is applied to measure the net pairwise volatility spillover effect between different stock market indices. The net volatility spillover graph of the $24$ stock market indices is presented below. The nodes represent different indices and the numbers on the edges measure the net pairwise volatility spillover effect. Larger values indicate stronger volatility spillover effects while smaller values reflect weaker volatility transmission.
Figure 2: Net volatility spillover graph.
The baseline models include the HAR, VHAR, HAR-KS and GNNHAR models. Although the exact construction of some baseline models is not available, they are formulated according to the corresponding research papers. Notably, for the GNNHAR model, different values are tried for the hyperparameter $L$ and only the best performance is presented in each RV forecasting task. The relevant dataset and the code of the proposed models can be found at https://github.com/MikeZChi/GSPHAR.git.
@@ 4.3 Experiment Results
where $\text{MAE}_{i}$ is the MAE of the $i_{\text{th}}$ stock market index and $T_{\text{out}}$ denotes the duration of the out-of-sample dataset. $\widehat{\mathbf{v}}_{i,t}$ refers to the square root RV forecasting value of the $i_{\text{th}}$ index at time $t$ and $\mathbf{v}_{i,t}$ is the true square root RV value of the $i_{\text{th}}$ index at time $t$. The index-wise MAE results produced by each model with different forecasting horizon ($h=1,5,22$) are presented in Tables 1, 2 and 3.
@@
$h=1$ HAR VHAR HAR-KS GNNHAR GSPHAR
@@
$h=5$ HAR VHAR HAR-KS GNNHAR GSPHAR
@@
$h=22$ HAR VHAR HAR-KS GNNHAR GSPHAR
@@
According to the MAE results in Tables 1, 2 and 3, it is clear that the GSPHAR model improves RV forecasting results over all three forecast horizons. The GSPHAR model achieves the minimum MAE scores for $24$, $20$ and $21$ stock market indices in the short-term ($h=1$), mid-term ($h=5$) and long-term ($h=22$) RV forecasting tasks, respectively. Even when it does not achieve the best performance, the GSPHAR model still produces the second smallest MAE scores. These imply the consistent superiority of the GSPHAR model in different forecasting horizons. Besides, linear HAR models are inferior to the GSPHAR model and the GNNHAR model. This implies that the linearity is insufficient in capturing the complex volatility spillover patterns.
@@
Table 4: The MCS test results table of the short-term forecasting task with bold numbers representing that the corresponding models are included in the MCS at the $5\%$ significance level.
$h=1$ HAR VHAR HAR-KS GNNHAR GSPHAR
FCHI $\mathbf{0.107}$ 00 $0.044$ $\mathbf{0.182}$ $\mathbf{1}$
AEX $010.0$ 00 $0.001$ $\mathbf{1}$
BFX $0.002$ $0.002$ $\mathbf{0.336}$ $0.015$ $\mathbf{1}$
STOXX50E $\mathbf{0.059}$ $0.003$ $\mathbf{0.059}$ $\mathbf{0.059}$ $\mathbf{1}$
IBEX $030.0$ 00 $0.034$ $\mathbf{1}$
GDAXI $010.0$ 00 $0.001$ $\mathbf{1}$
AORD $0.024$ $\mathbf{0.059}$ $\mathbf{0.781}$ $0.035$ $\mathbf{1}$
FTSE $0.003$ $\mathbf{0.133}$ $\mathbf{0.168}$ $0.005$ $\mathbf{1}$
MXX $\mathbf{0.265}$ $0.001$ $\mathbf{0.128}$ $0.007$ $\mathbf{1}$
IXIC $\mathbf{0.064}$ 00 $0.014$ $\mathbf{0.074}$ $\mathbf{1}$
SSMI $050.0$ 00 $\mathbf{0.237}$ $\mathbf{1}$
SPX $030.0$ 00 $\mathbf{0.121}$ $\mathbf{1}$
RUT $0.019$ $\mathbf{0.615}$ $\mathbf{0.948}$ $\mathbf{0.479}$ $\mathbf{1}$
DJI 00 00 00 $\mathbf{0.115}$ $\mathbf{1}$
BSESN 00 00 00 00 $\mathbf{1}$
NSEI 00 00 00 00 $\mathbf{1}$
KS11 00 00 00 $0.005$ $\mathbf{1}$
BVSP 00 00 00 $0.042$ $\mathbf{1}$
HSI $0.009$ $0.001$ $0.009$ $0.009$ $\mathbf{1}$
KSE 00 00 00 $0.002$ $\mathbf{1}$
N225 00 00 00 00 $\mathbf{1}$
SSEC 00 00 00 00 $\mathbf{1}$
OSEAX 00 00 $\mathbf{0.380}$ 00 $\mathbf{1}$
GSPTSE 00 00 00 $\mathbf{0.731}$ $\mathbf{1}$
@@
Table 5: The MCS test results table of the mid-term forecasting task with bold numbers representing that the corresponding models are included in the MCS at the $5\%$ significance level.
$h=5$ HAR VHAR HAR-KS GNNHAR GSPHAR
FCHI 00 00 00 00 $\mathbf{1}$
AEX $040.0$ 00 $0.002$ $\mathbf{1}$
BFX $110.0$ $0.011$ $0.011$ $\mathbf{1}$
STOXX50E 00 00 00 $0.001$ $\mathbf{1}$
IBEX 00 00 00 00 $\mathbf{1}$
GDAXI 00 00 00 00 $\mathbf{1}$
AORD $\mathbf{0.086}$ $0.001$ $\mathbf{0.753}$ $\mathbf{1}$ $\mathbf{0.753}$
FTSE $040.0$ $0.047$ $\mathbf{0.442}$ $\mathbf{1}$
MXX $\mathbf{0.454}$ 00 $\mathbf{0.454}$ 00 $\mathbf{1}$
IXIC $330.0$ $\mathbf{0.075}$ $\mathbf{0.403}$ $\mathbf{1}$
SSMI $090.0$ 00 $\mathbf{0.087}$ $\mathbf{1}$
SPX $020.0$ 00 $\mathbf{1}$ $\mathbf{0.379}$
RUT $\mathbf{0.375}$ $0.007$ $\mathbf{0.778}$ $\mathbf{0.778}$ $\mathbf{1}$
DJI $050.0$ $0.005$ $\mathbf{1}$ $\mathbf{0.480}$
BSESN 00 00 00 00 $\mathbf{1}$
NSEI 00 00 00 00 $\mathbf{1}$
KS11 $0.003$ $0.003$ $0.003$ $0.003$ $\mathbf{1}$
BVSP 00 00 $0.002$ $\mathbf{1}$ $\mathbf{0.161}$
HSI $210.0$ $0.021$ $0.007$ $\mathbf{1}$
KSE $0.001$ $0.006$ $090.0$ $\mathbf{1}$
N225 00 00 $0.001$ $\mathbf{0.807}$ $\mathbf{1}$
SSEC 00 00 00 00 $\mathbf{1}$
OSEAX 00 00 00 $0.012$ $\mathbf{1}$
GSPTSE $200.0$ 00 $\mathbf{0.166}$ $\mathbf{1}$
@@
Table 6: The MCS test results table of the long-term forecasting task with bold numbers representing that the corresponding models are included in the MCS at the $5\%$ significance level.
$h=22$ HAR VHAR HAR-KS GNNHAR GSPHAR
FCHI $090.0$ 00 $\mathbf{0.051}$ $\mathbf{1}$
AEX $040.0$ $0.001$ $0.021$ $\mathbf{1}$
BFX $170.0$ $0.006$ $0.017$ $\mathbf{1}$
STOXX50E $010.0$ 00 $0.007$ $\mathbf{1}$
IBEX $070.0$ 00 $0.007$ $\mathbf{1}$
GDAXI 00 00 00 $0.008$ $\mathbf{1}$
AORD $\mathbf{0.052}$ $\mathbf{0.052}$ $\mathbf{0.052}$ $\mathbf{0.052}$ $\mathbf{1}$
FTSE $0.007$ $0.001$ $0.004$ $0.007$ $\mathbf{1}$
MXX $\mathbf{0.594}$ 00 $\mathbf{0.509}$ 00 $\mathbf{1}$
IXIC $0.039$ $0.018$ $0.039$ $0.039$ $\mathbf{1}$
SSMI $010.0$ 00 $\mathbf{1}$ $\mathbf{0.887}$
SPX $140.0$ $0.001$ $\mathbf{0.200}$ $\mathbf{1}$
RUT $\mathbf{0.087}$ $0.006$ $0.032$ $0.012$ $\mathbf{1}$
DJI $060.0$ $0.001$ $\mathbf{0.128}$ $\mathbf{1}$
BSESN 00 00 00 $\mathbf{0.389}$ $\mathbf{1}$
NSEI 00 00 00 $\mathbf{0.390}$ $\mathbf{1}$
KS11 $040.0$ $0.008$ $0.021$ $\mathbf{1}$
BVSP $\mathbf{0.071}$ 00 $0.041$ $\mathbf{0.071}$ $\mathbf{1}$
HSI $\mathbf{0.134}$ 00 $0.003$ $0.037$ $\mathbf{1}$
KSE $270.0$ $\mathbf{0.201}$ $\mathbf{0.218}$ $\mathbf{1}$
N225 00 $050.0$ $\mathbf{0.175}$ $\mathbf{1}$
SSEC 00 00 00 $\mathbf{1}$ $\mathbf{0.570}$
OSEAX $050.0$ 00 $0.011$ $\mathbf{1}$
GSPTSE $030.0$ 00 $\mathbf{1}$ $\mathbf{0.920}$
@@
The MAE results and the MCS test demonstrate that the GSPHAR model improved the RV forecasting performance compared to the other tested models. To further evaluate the RV forecasting results of the GSPHAR model across different forecasting horizons, the one-sided DM test is applied to perform pairwise comparisons against the competing models. Suppose the two candidate models are model 00 and model $1$. The null hypothesis of the one-sided DM test is that model 00 and model $1$ are indifferent or model $1$ is inferior to model 00 in terms of forecast accuracy. When the DM test statistic is sufficiently large, the DM test can reject the null hypothesis and conclude that model $1$ provides a significantly more accurate forecast than model 00. The results of the DM test are presented in Tables 7, 8 and 9. Here, for each DM test model $1$ represents the GSPHAR model, and model 00 represents other models.
Table 7: The DM test results table of the short-term forecasting task with bold numbers represents that the GSPHAR is significantly better than the competing model at the $5\%$ significance level.